OpenStudy (anonymous):

How do you graph (y-3)=-3/4(x-6)?

OpenStudy (dtan5457):

That is currently in point slope form, and you should simplify it. Do you know how?

OpenStudy (anonymous):

Kinda

OpenStudy (dtan5457):

Remove the parenthesis (y-3)=-3/4(x-6) y-3=-3/4x+4.5 +3 +3 y=-3/4x+7.5

OpenStudy (dtan5457):

y=-3/4x+7.5 ^ that is now in slope intercept form. You can either use a graphing calculator, or make a table by hand and plug random x coordinates to get your outputs.

OpenStudy (dtan5457):

From y=-3/4x+7.5 You can tell the slope is -3/4 and the y intercept is 7.5

OpenStudy (anonymous):

Ohhh ok now i get it thanks!

OpenStudy (dtan5457):

No problem. Feel free to fan and medal. Tag me in anymore questions if needed :)